# Towards LLM Unlearning Resilient to Relearning Attacks: A Sharpness-Aware Minimization Perspective and Beyond

### v3

@@ 1 Introduction
Despite the increasing importance of LLM unlearning, several recent studies (Łucki et al., 2024; Zhang et al., 2024e; Lynch et al., 2024; Hu et al., 2024; Deeb & Roger, 2024) have identified a critical issue: LLM unlearning often lacks robustness. Specifically, the susceptibility to quickly recovering ‘already-unlearned’ knowledge post-unlearning is evident through so-called relearning attacks (Lynch et al., 2024; Hu et al., 2024). These attacks can effectively reverse the unlearning process by leveraging lightweight fine-tuning on the unlearned model using only a small number of data from the forget dataset.
Although numerous LLM unlearning methods have been proposed in the literature (Yao et al., 2024; Maini et al., 2024; Ji et al., 2024b; Zhang et al., 2024a; Liu et al., 2024a; Ji et al., 2024b; Li et al., 2024; Jia et al., 2024a; b), few studies have explored the robust optimization foundation for LLM unlearning. For example, negative preference optimization (NPO) (Zhang et al., 2024a), one of the state-of-the-art (SOTA) LLM unlearning methods, has demonstrated superior unlearning effectiveness compared to other approaches (Shi et al., 2024). However, as we will motivate in Sec. 3, NPO still remains vulnerable to relearning attacks. This highlights the need to develop a robust optimization foundation to strengthen LLM unlearning against such attacks. Tracing back to defenses against classic (input-level) prediction-evasion adversarial attacks, adversarial training (Madry et al., 2018), built upon min-max optimization, has proven to be a generic and effective robust optimization framework. In a similar vein, we ask:
@@ 2 Related Work
Machine unlearning modifies models to remove the influence of undesirable data, originally developed to mitigate post-training privacy risks (Cao & Yang, 2015; Ginart et al., 2019; Ullah et al., 2021; Fan et al., 2024c). While retraining from scratch guarantees exact unlearning, it is computationally prohibitive, leading to research on approximate unlearning methods that balance efficiency and effectiveness (Kurmanji et al., 2024; Fan et al., 2024a; Chen et al., 2023; Zhang et al., 2024c; d). A rapidly growing subfield is LLM unlearning (Jang et al., 2022; Meng et al., 2022; Yao et al., 2024; Eldan & Russinovich, 2023; Jia et al., 2024b; Zhang et al., 2024a; Maini et al., 2024; Jia et al., 2024a; Liu et al., 2024c; Fan et al., 2024b; Thaker et al., 2024), which has been shown promise in mitigating the generation of harmful content (Yao et al., 2024; Li et al., 2024; Jia et al., 2024b) and protecting sensitive, copyrighted, or private information (Eldan & Russinovich, 2023; Wu et al., 2023; Jang et al., 2022). Existing LLM unlearning approaches include model-based optimization (Maini et al., 2024; Yao et al., 2024; Jia et al., 2024a; Fan et al., 2024b; Zhang et al., 2024a; Li et al., 2024; Jia et al., 2024a; Wu et al., 2023; Fan et al., 2024b) and input-based strategies (via prompting or in-context learning) to facilitate unlearning without extensive parameter adjustments (Liu et al., 2024a; Thaker et al., 2024; Pawelczyk et al., 2023). Furthermore, recent benchmarking efforts provide valuable frameworks for evaluating the effectiveness of LLM unlearning approaches. These include TOFU (Maini et al., 2024), which focuses on fictitious unlearning using synthetic data, WMDP (Li et al., 2024), which aims to mitigate sociotechnical harms in model generation, and MUSE (Shi et al., 2024), which focuses on erasing copyrighted information from LLMs.
‘Adversaries’ in LLM unlearning.
@@
Recent studies have also exposed critical robustness vulnerabilities in existing LLM unlearning approaches (Lynch et al., 2024; Łucki et al., 2024; Hu et al., 2024; Zhang et al., 2024e; Shumailov et al., 2024; Barez et al., 2025; Patil et al., 2024; Deeb & Roger, 2024; Lo et al., 2024). These vulnerabilities primarily fall into two categories: relearning attacks (Hu et al., 2024; Lynch et al., 2024; Deeb & Roger, 2024; Lo et al., 2024), where fine-tuning with even a small subset of forget samples can restore unlearned knowledge (Lynch et al., 2024), and jailbreaking attacks (Łucki et al., 2024; Lynch et al., 2024; Patil et al., 2024), where adversarial prompts successfully recover forgotten information at inference time (Łucki et al., 2024). To enhance the robustness of LLM unlearning, (Tamirisa et al., 2024) utilized a model-agnostic meta-learning (MAML) framework (Nichol, 2018) to counter tampering attacks, while (Sheshadri et al., 2024) employed adversarial training in the latent space of LLMs. Unlike existing work, we investigate unlearning robustness against relearning attacks through the lens of smoothness optimization, establishing a seamless connection to SAM, a direct yet underexplored optimization foundation for robust LLM unlearning.
SAM and smoothness optimization.
@@
Sharpness-aware minimization (SAM) is a representative smoothness optimization technique that minimizes both the loss value and its sharpness, effectively promoting a flatter loss landscape, originally introduced to improve model generalization (Foret et al., 2021; Andriushchenko & Flammarion, 2022; Liu et al., 2022b; Du et al., 2022; Zhang et al., 2023; Jiang et al., 2023). SAM has also been applied in traditional adversarial training to defend against input-level adversarial attacks (Wei et al., 2023; Zhang et al., 2024b). Beyond SAM, other smoothness optimization approaches include gradient penalty (GP) and curvature regularization (CR), which impose penalties based on loss gradients or Hessian-gradient products to encourage smoothness (Dauphin et al., 2024; Zhao et al., 2024). Randomized smoothing (RS) improves smoothness by convolving a non-smooth objective function with a Gaussian distribution (Duchi et al., 2012; Cohen et al., 2019; Ji et al., 2024a). Meanwhile, weight averaging (WA) enhances smoothness by averaging model weights across training iterations, leading to a smoother optimization trajectory (Izmailov et al., 2018). Further, activation smoothing has been used to mitigate the security risk of users fine-tuning for LLMs (Huang et al., 2024). These smoothness optimization approaches will serve as a key foundation for enhancing the robustness of LLM unlearning in this work.
@@ 3 LLM Unlearning and Relearning Attacks
Preliminaries on unlearning and relearning attacks.
To achieve efficient LLM unlearning while preserving model utility, the unlearning problem is formulated as an optimization task to update parameters from their pretrained values (Eldan & Russinovich, 2023; Yao et al., 2024; Maini et al., 2024; Zhang et al., 2024a; Li et al., 2024). To be specific, let $\mathcal{D}_{\mathrm{f}}$ and $\mathcal{D}_{\mathrm{r}}$ represent the ‘forget’ and ‘retain’ sets, respectively. Here the forget set $\mathcal{D}_{\mathrm{f}}$ defines the scope of unlearning, specifying the data samples whose influences are to be removed. Conversely, the retain set $\mathcal{D}_{\mathrm{r}}$ ensures the preservation of the model’s utility post-unlearning. Built upon $\mathcal{D}_{\mathrm{f}}$ and $\mathcal{D}_{\mathrm{r}}$, a forget loss ($\ell_{\mathrm{f}}$) and a retain loss ($\ell_{\mathrm{r}}$) are defined to balance unlearning effectiveness and utility retention. The leads to the following regularized optimization problem (Liu et al., 2024b):
@@
A motivating example.
Fig. 1 presents the performance of the NPO-based unlearning approach to solve (2) in mitigating the malicious use of the LLM Zephyr-7B-beta on the WMDP (Weapons of Mass Destruction Proxy) Bio dataset (Li et al., 2024). In this context, a lower accuracy of the model on the WMDP (Bio) evaluation set corresponds to better unlearning. Thus, we define unlearning effectiveness (UE) as 1-Accuracy on WMDP evaluation set, where a higher value indicates better unlearning performance.
@@
Sharpness-aware minimization (SAM): A robust optimization perspective on unlearning against relearning.
Building on (2) and (4), enhancing unlearning resistance to relearning attacks can be framed as an adversary-defense game. This framework, similar to adversarial training (Madry et al., 2018), can be expressed using min-max optimization, where the objective is to jointly optimize the unlearning process to counteract the adversarial relearning attempts effectively. However, unlike adversarial training, which defends against input-level adversarial examples, relearning attacks directly modify the weights of the unlearned model to counteract the forget objective. If the relearning objective $\ell_{\mathrm{relearn}}$ is defined to counteract the forget objective, such that $\ell_{\mathrm{relearn}}=-\ell_{\mathrm{f}}$, then integrating the relearning adversary (4) into LLM unlearning (2) leads to the following min-max robust optimization problem:
@@ 4 Enhancing Unlearning Robustness: From SAM to Broader Smoothness Optimization
SAM facilitates curvature regularization of forget loss.
As shown by (6), SAM promotes the flatness of the forget loss landscape since it seeks a minimum that maintains a uniformly low loss across the neighborhood of the model. Therefore, SAM facilitates smoothness optimization in LLM unlearning. Based on the SAM algorithm (Foret et al., 2021), the inner maximization in (6) can be solved in closed form using linear approximation:
$\bm{\delta}^{*}({\bm{\theta}})\mathrel{\mathop{:}}=$ $\displaystyle\operatorname*{arg\,max}_{\|\bm{\delta}\|_{2}\leq\rho}\ell_{\mathrm{f}}({\bm{\theta}}+\bm{\delta})\overset{(a)}{\approx}\displaystyle\operatorname*{arg\,max}_{\|\bm{\delta}\|_{2}\leq\rho}\ell_{\mathrm{f}}({\bm{\theta}})+\bm{\delta}^{\top}\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})=\displaystyle\operatorname*{arg\,max}_{\|\bm{\delta}\|_{2}\leq\rho}\bm{\delta}^{\top}\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\overset{(b)}{=}\rho\frac{\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})}{\|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\|_{2}},$ (7)
@@
$\ell^{\mathrm{SAM}}_{\mathrm{f}}({\bm{\theta}})=\ell_{\mathrm{f}}\left({\bm{\theta}}+\rho\frac{\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})}{\|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\|_{2}}\right)\approx\ell_{\mathrm{f}}({\bm{\theta}})+\rho\frac{\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})^{\top}\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})}{\|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\|_{2}}=\ell_{\mathrm{f}}({\bm{\theta}})+\rho\|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\|_{2}.$ (9)
@@
$\frac{d\|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\|_{2}}{d{\bm{\theta}}}=\frac{d(\|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\|_{2}^{2})^{1/2}}{d{\bm{\theta}}}=\frac{1}{2}(\|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\|_{2}^{2})^{-1/2}(2\mathbf{H}\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}}))=\mathbf{H}\mathbf{v},$ (10)
@@
Broader smoothness optimization to improve unlearning robustness.
As analyzed above, the SAM-like optimization in (6) and (8) indicates smoothness optimization for robust unlearning against relearning attacks. Building on this insight, we extend our investigation to a broader range of smoothness optimization techniques, including randomized smoothing (RS), gradient penalty (GP), curvature regularization (CR), and weight averaging (WA).
@@
Smoothness in unlearning improves robustness: A loss landscape perspective.
Furthermore, we investigate the previously discussed smoothness optimization techniques (SAM, RS, GP, CR, and WA) and their role in enhancing unlearning robustness, through the perspective of the loss landscape. The loss landscape represents the geometric surface of a loss function against its model parameter change (Li et al., 2018; Hao et al., 2019; Zan et al., 2022). For ease of visualization, the loss sensitivity can be assessed using a parametric model defined as $f(x,y)=\ell({\bm{\theta}}+x\cdot\mathbf{r}_{1}+y\cdot\mathbf{r}_{2})$. Here, $\ell$ represents the prediction loss function, $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are two directional vectors given by Gaussian vectors, and $x$ and $y$ are scalar parameters that define the perturbation strength. The 3D loss landscape visualization is subsequently achieved by plotting the loss sensitivity w.r.t. the perturbation parameters $x$ and $y$. Smoothness is indicated when the loss landscape appears relatively flat in the vicinity of the current model parameters.
@@
Figs. 2-(b)$\sim$(c) illustrate the prediction loss landscape of the original model and the NPO-unlearned model evaluated on the forget set $\mathcal{D}_{\mathrm{f}}$. The prediction loss is given by the cross-entropy next-token prediction loss. The $z$-axis represents the prediction loss, where higher values indicate more effective unlearning (i.e., worse prediction performance). As observed, NPO increases the prediction loss on $\mathcal{D_{\mathrm{f}}}$ at $x=y=0$, indicating effective unlearning. Without the application of smoothness-promoting techniques, the vanilla loss landscape is notably sharp around $x=y=0$, corresponding to the neighborhood of the unlearned model. In contrast, Figs. 2-(d)$\sim$(h) depict the loss landscapes of unlearned models employing the smooth variants of NPO. As we can see, the loss landscape becomes significantly smoother than Fig. 2-(c) when using SAM, RS, GP, CR, and WA. Taken together, Fig. 2 shows that the smoothness of the loss landscape is beneficial to unlearning robustness improvement. We also provide the loss landscape on $\mathcal{D}_{\mathrm{r}}$ in Figs. A1 of Appendix B for comparison.

### v4

@@ 1 Introduction
Despite the increasing importance of LLM unlearning, several recent studies (Łucki et al., 2024; Zhang et al., 2024c; Lynch et al., 2024; Hu et al., 2024; Deeb & Roger, 2024) have identified a critical issue: LLM unlearning often lacks robustness. Specifically, the susceptibility to quickly recovering ‘already-unlearned’ knowledge post-unlearning is evident through so-called relearning attacks (Lynch et al., 2024; Hu et al., 2024). These attacks can effectively reverse the unlearning process by leveraging lightweight fine-tuning on the unlearned model using only a small number of data from the forget dataset.
Although numerous LLM unlearning methods have been proposed in the literature (Yao et al., 2024; Maini et al., 2024; Ji et al., 2024b; Zhang et al., 2024a; Liu et al., 2024a; Ji et al., 2024b; Li et al., 2024; Jia et al., 2024a, b), few studies have explored the robust optimization foundation for LLM unlearning. For example, negative preference optimization (NPO) (Zhang et al., 2024a), one of the state-of-the-art (SOTA) LLM unlearning methods, has demonstrated superior unlearning effectiveness compared to other approaches (Shi et al., 2024). However, as we will motivate in Sec. 3, NPO still remains vulnerable to relearning attacks. This highlights the need to develop a robust optimization foundation to strengthen LLM unlearning against such attacks. Tracing back to defenses against classic (input-level) prediction-evasion adversarial attacks, adversarial training (Madry et al., 2018), built upon min-max optimization, has proven to be a generic and effective robust optimization framework. In a similar vein, we ask:
@@ 2 Related Work
Machine unlearning and its applications to LLMs. Machine unlearning modifies models to remove the influence of undesirable data, originally developed to mitigate post-training privacy risks (Cao & Yang, 2015; Ginart et al., 2019; Ullah et al., 2021). While retraining from scratch guarantees exact unlearning, it is computationally prohibitive, leading to research on approximate unlearning methods that balance efficiency and effectiveness (Kurmanji et al., 2024; Fan et al., 2024b; Chen et al., 2023). A rapidly growing subfield is LLM unlearning (Jang et al., 2022; Meng et al., 2022; Yao et al., 2023; Eldan & Russinovich, 2023; Jia et al., 2024b; Zhang et al., 2024a; Maini et al., 2024; Jia et al., 2024a; Liu et al., 2024c; Fan et al., 2024a; Thaker et al., 2024), which has been shown promise in mitigating the generation of harmful content (Yao et al., 2023; Li et al., 2024; Jia et al., 2024b) and protecting sensitive, copyrighted, or private information (Eldan & Russinovich, 2023; Wu et al., 2023; Jang et al., 2022). Existing LLM unlearning approaches include model-based optimization (Maini et al., 2024; Yao et al., 2023; Jia et al., 2024a; Fan et al., 2024a; Zhang et al., 2024a; Li et al., 2024; Jia et al., 2024a; Wu et al., 2023; Fan et al., 2024a) and input-based strategies (via prompting or in-context learning) to facilitate unlearning without extensive parameter adjustments (Liu et al., 2024a; Thaker et al., 2024; Pawelczyk et al., 2023). Furthermore, recent benchmarking efforts provide valuable frameworks for evaluating the effectiveness of LLM unlearning approaches. These include TOFU (Maini et al., 2024), which focuses on fictitious unlearning using synthetic data, WMDP (Li et al., 2024), which aims to mitigate sociotechnical harms in model generation, and MUSE (Shi et al., 2024), which focuses on erasing copyrighted information from LLMs.
@@
‘Adversaries’ in LLM unlearning. Recent studies have also exposed critical robustness vulnerabilities in existing LLM unlearning approaches (Lynch et al., 2024; Łucki et al., 2024; Hu et al., 2024; Zhang et al., 2024c; Shumailov et al., 2024; Barez et al., 2025; Patil et al., 2024; Deeb & Roger, 2024). These vulnerabilities primarily fall into two categories: relearning attacks (Hu et al., 2024; Lynch et al., 2024; Deeb & Roger, 2024), where fine-tuning with even a small subset of forget samples can restore unlearned knowledge (Lynch et al., 2024), and jailbreaking attacks (Łucki et al., 2024; Lynch et al., 2024; Patil et al., 2024), where adversarial prompts successfully recover forgotten information at inference time (Łucki et al., 2024). To enhance the robustness of LLM unlearning, Tamirisa et al. (2024) utilized a model-agnostic meta-learning (MAML) framework (Nichol, 2018) to counter tampering attacks, while Sheshadri et al. (2024) employed adversarial training in the latent space of LLMs. Unlike existing work, we investigate unlearning robustness against relearning attacks through the lens of smoothness optimization, establishing a seamless connection to SAM, a direct yet underexplored optimization foundation for robust LLM unlearning.
@@
SAM and smoothness optimization. Sharpness-aware minimization (SAM) is a representative smoothness optimization technique that minimizes both the loss value and its sharpness, effectively promoting a flatter loss landscape, originally introduced to improve model generalization (Foret et al., 2021; Andriushchenko & Flammarion, 2022; Liu et al., 2022b; Du et al., 2022; Zhang et al., 2023). SAM has also been applied in traditional adversarial training to defend against input-level adversarial attacks (Wei et al., 2023; Zhang et al., 2024b). Beyond SAM, other smoothness optimization approaches include gradient penalty (GP) and curvature regularization (CR), which impose penalties based on loss gradients or Hessian-gradient products to encourage smoothness (Dauphin et al., 2024; Zhao et al., 2024). Randomized smoothing (RS) improves smoothness by convolving a non-smooth objective function with a Gaussian distribution (Duchi et al., 2012; Cohen et al., 2019; Ji et al., 2024a). Meanwhile, weight averaging (WA) enhances smoothness by averaging model weights across training iterations, leading to a smoother optimization trajectory (Izmailov et al., 2018). These smoothness optimization approaches will serve as a key foundation for enhancing the robustness of LLM unlearning in this work.
@@ 3 LLM Unlearning and Relearning Attacks
Preliminaries on unlearning and relearning attacks. To achieve efficient LLM unlearning while preserving model utility, the unlearning problem is formulated as an optimization task to update parameters from their pretrained values (Eldan & Russinovich, 2023; Yao et al., 2024; Maini et al., 2024; Zhang et al., 2024a; Li et al., 2024). To be specific, let $\mathcal{D}_{\mathrm{f}}$ and $\mathcal{D}_{\mathrm{r}}$ represent the ‘forget’ and ‘retain’ sets, respectively. Here the forget set $\mathcal{D}_{\mathrm{f}}$ defines the scope of unlearning, specifying the data samples whose influences are to be removed. Conversely, the retain set $\mathcal{D}_{\mathrm{r}}$ ensures the preservation of the model’s utility post-unlearning. Built upon $\mathcal{D}_{\mathrm{f}}$ and $\mathcal{D}_{\mathrm{r}}$, a forget loss ($\ell_{\mathrm{f}}$) and a retain loss ($\ell_{\mathrm{r}}$) are defined to balance unlearning effectiveness and utility retention. The leads to the following regularized optimization problem (Liu et al., 2024b):
@@
A motivating example. Fig. 1 presents the performance of the NPO-based unlearning approach to solve (2) in mitigating the malicious use of the LLM Zephyr-7B-beta on the WMDP (Weapons of Mass Destruction Proxy) Bio dataset (Li et al., 2024). In this context, a lower accuracy of the model on the WMDP (Bio) evaluation set corresponds to better unlearning. Thus, we define unlearning effectiveness (UE) as 1-Accuracy on WMDP evaluation set, where a higher value indicates better unlearning performance.
@@
Sharpness-aware minimization (SAM): A robust optimization perspective on unlearning against relearning. Building on (2) and (4), enhancing unlearning resistance to relearning attacks can be framed as an adversary-defense game. This framework, similar to adversarial training (Madry et al., 2018), can be expressed using min-max optimization, where the objective is to jointly optimize the unlearning process to counteract the adversarial relearning attempts effectively. However, unlike adversarial training, which defends against input-level adversarial examples, relearning attacks directly modify the weights of the unlearned model to counteract the forget objective. If the relearning objective $\ell_{\mathrm{relearn}}$ is defined to counteract the forget objective, such that $\ell_{\mathrm{relearn}}=-\ell_{\mathrm{f}}$, then integrating the relearning adversary (4) into LLM unlearning (2) leads to the following min-max robust optimization problem:
@@ 4 Enhancing Unlearning Robustness: From SAM to Broader Smoothness Optimization
SAM facilitates curvature regularization of forget loss. As shown by (6), SAM promotes the flatness of the forget loss landscape since it seeks a minimum that maintains a uniformly low loss across the neighborhood of the model. Therefore, SAM facilitates smoothness optimization in LLM unlearning. Based on the SAM algorithm (Foret et al., 2021), the inner maximization in (6) can be solved in closed form using linear approximation:
$\bm{\delta}^{*}({\bm{\theta}})\mathrel{\mathop{:}}=$ $\displaystyle\operatorname*{arg\,max}_{\|\bm{\delta}\|_{2}\leq\rho}\ell_{\mathrm{f}}({\bm{\theta}}+\bm{\delta})\overset{(a)}{\approx}\displaystyle\operatorname*{arg\,max}_{\|\bm{\delta}\|_{2}\leq\rho}\ell_{\mathrm{f}}({\bm{\theta}})+\bm{\delta}^{\top}\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})$
$=$ $\displaystyle\operatorname*{arg\,max}_{\|\bm{\delta}\|_{2}\leq\rho}\bm{\delta}^{\top}\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\overset{(b)}{=}\rho\frac{\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})}{\|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\|_{2}},$ (7)
@@
$\ell^{\mathrm{SAM}}_{\mathrm{f}}({\bm{\theta}})=\ell_{\mathrm{f}}\left({\bm{\theta}}+\rho\frac{\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})}{\|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\|_{2}}\right)$
$\approx$ $\ell_{\mathrm{f}}({\bm{\theta}})+\rho\frac{\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})^{\top}\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})}{\|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\|_{2}}=\ell_{\mathrm{f}}({\bm{\theta}})+\rho\|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\|_{2}.$ (9)
@@
$\frac{d\|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\|_{2}}{d{\bm{\theta}}}=\frac{d(\|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\|_{2}^{2})^{1/2}}{d{\bm{\theta}}}$
$=$ $\frac{1}{2}(\|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}})\|_{2}^{2})^{-1/2}(2\mathbf{H}\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}}))=\mathbf{H}\mathbf{v},$ (10)
@@
Broader smoothness optimization to improve unlearning robustness. As analyzed above, the SAM-like optimization in (6) and (8) indicates smoothness optimization for robust unlearning against relearning attacks. Building on this insight, we extend our investigation to a broader range of smoothness optimization techniques, including randomized smoothing (RS), gradient penalty (GP), curvature regularization (CR), and weight averaging (WA).
@@
Smoothness in unlearning improves robustness: A loss landscape perspective. Furthermore, we investigate the previously discussed smoothness optimization techniques (SAM, RS, GP, CR, and WA) and their role in enhancing unlearning robustness, through the perspective of the loss landscape. The loss landscape represents the geometric surface of a loss function against its model parameter change (Li et al., 2018; Hao et al., 2019; Zan et al., 2022). For ease of visualization, the loss sensitivity can be assessed using a parametric model defined as $f(x,y)=\ell({\bm{\theta}}+x\cdot\mathbf{r}_{1}+y\cdot\mathbf{r}_{2})$. Here, $\ell$ represents the prediction loss function, $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are two directional vectors given by Gaussian vectors, and $x$ and $y$ are scalar parameters that define the perturbation strength. The 3D loss landscape visualization is subsequently achieved by plotting the loss sensitivity w.r.t. the perturbation parameters $x$ and $y$. Smoothness is indicated when the loss landscape appears relatively flat in the vicinity of the current model parameters.
@@
Figs. 2-(b)$\sim$(c) illustrate the prediction loss landscape of the original model and the NPO-unlearned model evaluated on the forget set $\mathcal{D}_{\mathrm{f}}$. The prediction loss is defined as $p_{\bm{\theta}}(y|x)=\frac{1}{|y|}\sum_{i=1}^{|y|}\log\pi_{\bm{\theta}}(y_{i}|x,y_{<i})$. The $z$-axis represents the prediction loss, where higher values indicate more effective unlearning (i.e., worse prediction performance). As observed, NPO increases the prediction loss on $\mathcal{D_{\mathrm{f}}}$ at $x=y=0$, indicating effective unlearning. Without the application of smoothness-promoting techniques, the vanilla loss landscape is notably sharp around $x=y=0$, corresponding to the neighborhood of the unlearned model. In contrast, Figs. 2-(d)$\sim$(h) depict the loss landscapes of unlearned models employing the smooth variants of NPO. As we can see, the loss landscape becomes significantly smoother than Fig. 2-(c) when using SAM, RS, GP, CR, and WA. Taken together, Fig. 2 shows that the smoothness of the loss landscape is beneficial to unlearning robustness improvement. We also provide the loss landscape on $\mathcal{D}_{\mathrm{r}}$ in Figs. A1 of Appendix B for comparison.